OpenStudy (anonymous):

HELP WIL FAN AND MEDAL! Solve for n. Round to the nearest tenth, if necessary 47/n=36/54

OpenStudy (anonymous):

The options are: A) n=3.1 B) n=70.5 C) n=0.7 D) n=2538

OpenStudy (rainbow_rocks03):

What do you think the answer is?

OpenStudy (anonymous):

I'm not sure.

OpenStudy (anonymous):

Could someone at leased show me how to set up the problem?

OpenStudy (rainbow_rocks03):

here is a note in rounding...5 or more raise the score,4 or less let it rest

OpenStudy (texaschic101):

or set it up as a proportion 47/n = 36/54 --- cross multiply (36)(n) = (47)(54) 36n = 2538 n = 2538/36 n = you can do the math

OpenStudy (anonymous):

Thank you guys I just figured it out!